OpenStudy (anonymous):

2x -(3x + 1) + 2x =5 Broken down?

OpenStudy (anonymous):

\[2x - (3x+1) + 2x = 5 \] \[2x - 3x - 1 + 2x = 5\] Do you need more help after this?

OpenStudy (anonymous):

if you can

OpenStudy (anonymous):

Do you know what \(2x- 2x\) is?

OpenStudy (anonymous):

it cancels out right?

OpenStudy (anonymous):

Yes and \(2x +2x\) is?

OpenStudy (anonymous):

4x obviously haha

OpenStudy (anonymous):

Then we reorder the question slightly \[2x+2x-3x-1 = 5\] How many x do we have?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

well after you add the 2x + 2x you get 4x - 3x so there would be 2 x's amd then it would just end up being "x" correct??

OpenStudy (anonymous):

Yes, so now you have \[x-1=5\] Do yo know how to solve it now?

OpenStudy (anonymous):

What does x have to be so that if you subtract 1 from it it becomes 5?

OpenStudy (anonymous):

add 1 to 5 right?

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

And there you have it :)

OpenStudy (anonymous):

thank you :) soooooooooo much!

OpenStudy (anonymous):

Np, you solved it yourself you know

OpenStudy (anonymous):

well yes but with your guidance!!